OpenStudy (anonymous):

A culture started with 5,000 bacteria. After 3 hours, it grew to 6,500 bacteria. How many bacteria will be present after 13 hours. P=Ae^kt

OpenStudy (anonymous):

create a proportion 3/6500=13/x 13*6500= 84500 84500/3= 28166.7 hope this helped

OpenStudy (asnaseer):

this is not a linear equation so you cannot use proportions here

OpenStudy (asnaseer):